Review 1 Constants Part A A spring has an unstretched length of 14 cm. When an 84 g ball is hung from it, the length increases by 7.0 cm. Then the ball is pulled down another 7.0 cm and released What is the spring constant of the spring? Express your answer using two significant figures k= N/m Submit Part B What is the period of the oscillation? Express your answer using two significant figures T- Submit Request Answer
Answer #1

Part A

We know that

Magnitude of the spring force

F=Ky. ....1

Here F=mg . m=84 g=84×10^(-3) kg , g=9.81 m/s^2 ,y= 7cm

y=7×10^(-2)m

So from equation ...1

K= mg/y =84×10^(-3)×9.81/7×10^(-2)

K=11.772 N/m

K=11.8 N/m. Answer

Part B

Period of the spring .

T^2= 4×π^2×m/K

T^2=4×π^2×84×10^(-3)/11.8

T^2=0.281

T=0.53 s answer